# ChemKANs for Combustion Chemistry Modeling and Acceleration

### v1

@@ 2.3.1 Vanilla KAN-ODEs
where each is a unique learnable activation function connecting a single input to a single output (thus ). In other words, each input is connected to each output with a unique learnable activation function (much like in an MLP, where each input is connected to each output with a unique learnable weight), leading to a total of activation functions connecting the and layers. We use RBF basis functions in the current work as in [32, 6, 33], although the choice of is is flexible, and many other options have been proposed in the literature including B-splines [23], ReLU functions [34], and various other combinations [35, 36]. The AddKAN structure has an inherent problem of expressing the equation only with additive operations, limiting its concise expressivity for multiplicative operations. Therefore, recent studies have proposed new layer structures to address this issue and improve parameter efficiency [37, 33]. Here we introduce LeanKAN [33] which has shown promise to be the most effective and efficient for both additive and multiplicative operations. The LeanKAN structure is achieved by summation of two separate terms, and , such that
where is the multiplication hyperparameter that dictates the number of multiplication input nodes, and are the same univariate activation functions used in AddKAN (to be defined below). Overall, the LeanKAN formulation takes the matrix of activated inputs, computes sums and products in parallel according to to construct the intermediate , and finally sums the multiplication and addition components of to arrive at the next layer. Further derivation and implementation details are available for KANs in [23], KAN-ODEs in [27], and LeanKAN layers in [33].
@@
where we define as the grid size, or number of superimposed basis functions used to construct a single activation. and are the learnable network parameters that make up , where scales each gridpoint’s RBF basis functions within the sum, and scales a single base activation function . and are the index of the layer ’s matrix in Eq. 7 to which the weights are applied, while the grid itself is defined by its individual gridpoints and gridpoint spacing (or RBF spreading parameter) . While the original work of [23] suggests periodically re-gridding these basis functions to ensure they are learning on the proper input range, here we instead normalize at each layer input as in more recent works [38, 39, 27, 33] for computational efficiency. The single base activation term in each is a Swish activation function [40].
2.3.2 Novel chemistry KAN (ChemKAN) architecture
@@
We define the terms in these equations line by line throughout this paragraph. In Eq. 13, we encode prior knowledge of the true species production/consumption rate relationship of Eq. 1 into the ChemKAN by computing the species-only production rate from the entire state input , via a KAN network (KAN) parameterized by . Then, from the true energy equation (Eq. 2), we recognize that the temperature rate is a simple linear combination of the species production rates, with scaling factors defined by the enthalpy and specific heat values (or alternatively, in Eq. 14 by the scalars in the linear mapping parameterized by ). Following from this, the crux of the thermodynamic superstructure is a computationally trivial, simple linear sum of the already-evaluated outputs of the kinetic core, as shown in Eqs. 14 and 15. One level deeper, we recognize that a secondary effect in the true Eq. 2 is the dependence of the thermophysical parameters, specifically , on the temperature and species mixture. The error stemming from the first-order Linear() approximation’s failure to account for such thermophysical parameter variation is reflected in the term of Eq. 14, which in the final formulation of Eq. 15 is accounted for via a supplemental single-layer, single-output KAN correction carrying forward a functional dependence on the species and temperature inputs, parameterized by . Overall, ChemKAN is composed of a kinetic core structure and a thermodynamic superstructure that strongly mimic the true governing equations, operate largely in series, and include full sharing of all reaction and species production information. This architecture allows for versatile and flexible modeling by turning “on” or “off” the energy equation for standalone, kinetic core-only modeling or combined kinetic and thermodynamic modeling.
2.3.3 Network structure details
The ChemKAN structure proposed here is combined with the KAN-ODE framework detailed in Sec. 2.3.1 to leverage both the physics-guided ChemKAN architecture and the proven dynamical system benefits offered by the KAN-ODE integration. For nomenclature consistency, we henceforth refer to both the newly designed network structure as well as the complete integration with an ODE solver as ChemKAN. For the relatively larger kinetic core responsible for learning all reaction information, we define the internal layer structure as
a two-layer structure as per the notation introduced in Eq. 6. The first is an AddKAN layer [23] (which is equivalently the LeanKAN with [33]). The second is a LeanKAN layer with to inject the multiplication operator. Further discussion of this specific two-layer form and its merits are provided in detail in Ref. [33]. The thermodynamic correction structure simply comprises a sum of univariate activations evaluated on the components of ,
@@ 2.3.4 Loss function
Here denotes thermochemical states normalized to the [0, 1] window by subtracting the minimum then dividing by the range, while and , respectively, are the network prediction and training data. In the first MSE term, when training the kinetic core, as only species profiles are learned. For the thermodynamic superstructure, is used to train the added temperature output. We also provide an optional element conservation physics-informed loss term, or PINN term [41], similar to Ref. [42] to encourage the ChemKAN to find models that obey physical laws. There, is the number of elements in the data (i.e., H, O, N), is the number of datapoints in the temporal profiles and is the number of species represented in the data. For a given element , the element conservation term computes the mass fraction difference across all species between the training data and ChemKAN reconstruction (). This is then converted to an elemental difference via , the atom count of element in species ; , the atomic mass of element ; and , the molar mass of species . This is computed across all elements , and finally weighted by (here ) before summing with the standard MSE term. In the examples below, we use the MSE loss only for all biodiesel model inference results, and MSE loss with the PINN term for H2 model acceleration.
2.3.5 Training Process
We highlight again that this architecture incorporates significant inductive bias by separating the inference of the kinetic behavior and thermodynamic behavior into distinct layers and activations within the complete structure, giving the network an explicit functional coupling between the temperature rate and the species rates to ease its training burden while retaining an accurate model of the real exothermic behavior of the system. In itself, this serially-run inductive split should theoretically ease some of the nonlinear training difficulties observed in [17] while retaining a single cohesive and information-sharing structure. Additionally, and more importantly to stable training and convergence behavior, it allows us to separate the training itself into a kinetic and a thermodynamic stage to greatly reduce the inference burden, as discussed here:
1.
Training stage 1—core kinetics : All inputs are used, and only the species production rate outputs are computed (see Eq. 13). The thermodynamic superstructure of Eqs. 14 and 15 is not used in this stage, and the input temperatures are simply read in from the training data to provide the kinetic core of the network with a simpler training task. See the grey highlight in Fig. 2. For cases without heat release, this step in isolation is sufficient for a complete model. With heat release, we move to stage 2 once stage 1 is converged.
2.
@@ 2.3.5 Training Process
This two-stage training process contrasts with the stage training process of ChemNODE [17], which requires networks to all learn their own distinct representations of what we know to be shared kinetic and thermodynamic governing laws. The current ChemKAN approach, while trained in two distinct stages, has an overall structure resembling a simple feedforward KAN thanks to its stacked design where all kinetic and thermodynamic information is shared across all inputs and outputs. For downstream evaluation, a single forward pass through the combined network of Fig. 2, or alternatively through Eqs. 1315, predicts the complete thermochemical state vector .
The learnable parameters are trained by the Adam optimizer [43]. While adjoint sensitivity analysis was used in the original KAN-ODE paper [27] to compute the gradients of a loss function , in the current work we note the numerical instability present in many chemical kinetic modeling problems due to numerical stiffness [44, 22, 18, 17, 3], potentially leading to failures with adjoint sensitivity analysis in the training process [3]. To prevent this, we implement forward sensitivity analysis to mitigate this potential stiffness issue in the ODE solver [3]. The learning rate was set to .
2.4 Data generation for case studies
We introduce two chemical reaction systems to demonstrate ChemKANs as a multi-purpose modeling technique. The first case of biodiesel production will illustrate the effectiveness of ChemKANs for inferring a chemical kinetic model purely from data. Then, we show computational acceleration for high-fidelity simulations using ChemKANs with an example of hydrogen-air combustion subject to significant numerical stiffness.
2.4.1 Model inference from noisy data – biodiesel production
@@
where the reaction rates scale with temperature via the standard Arrhenius law as , with for the three reactions. As in Ref. [5], we generate data in this case using kcal/mol and , with isothermal experiments at temperatures randomly sampled in the range of K to K. We define the species scalar quantities here as concentrations rather than mass fractions, to match convention with the governing equations. Initial TG and ROH concentrations are randomly sampled uniformly between and with all other intermediate and output species initialized at zero. 20 training data sets and 10 testing data sets are generated, with 30-second time windows in both consisting of sampled points. The temperature-dependent yet isothermal reaction rates present in this system motivate use of the kinetic ChemKAN core structure only. In other words, we drop the thermodynamic superstructure of Eq. 15.
@@
In this case we additionally probe the effectiveness of ChemKANs in the presence of significant experimental noise. To do so, we add increasing amounts of noise to the data and evaluate the ChemKAN’s performance against that of a standard Deep Operator Network [20].
2.4.2 Model acceleration – hydrogen-air combustion
@@
Despite its relatively simple chemistry, Hydrogen-air combustion is subject to stiffness, leading to high computational costs and numerical instability in the ODE solver. The main goal here is to achieve high accuracy in species and temperature profile reconstruction while reducing computational cost compared to the detailed chemistry solver using ChemKANs. We generate training data in Cantera [47], using the H2/O2 mechanism from GRI-Mech 3.0 [48] (9 species, 29 reactions). Data are generated at all 36 combinations of initial temperatures in {950, 1000, 1050, 1100, 1150, 1200} K and equivalence ratios in {0.7, 0.9, 1.1, 1.3, 1.5}, with the [1150 K, 1.3] case withheld as unseen testing data. The problem setup here is largely identical to that studied originally in ChemNODE [17] to facilitate fair comparison, although here we add the single withheld testing dataset to probe the ChemKAN’s robustness (thus 35 training datasets are used, compared to the 36 in ChemNODE).
@@
In addition to significantly increased numerical stiffness and complexity, a key fundamental difference between the current combustion system and the previous biodiesel synthesis case is the presence of substantial two-way temperature coupling as per Eqs. 12. To account for this, we include the thermodynamic ChemKAN superstructure and the two-stage training process outlined in Secs. 2.3.22.3.5.
3 Results and Discussion
3.1 Biodiesel model inference from noisy data
@@
We begin with analysis of the biodiesel model inference of the ChemKAN (the kinetic core only for this isothermal case), and a comparison against a traditional DeepONet for an identical task. DeepONets were selected as the target of comparison due to their general use in the scientific machine learning community, to provide a baseline against which to evaluate our first ChemKAN results. Note that comparison against ChemNODE is saved for the hydrogen-air combustion case in Sec. 3.2.
3.1.1 Model performance of the proposed ChemKAN architecture
@@
The kinetic core (Eq. 13) of the ChemKAN comprises two layers: one AddKAN and one LeanKAN, as discussed in Sec. 2.3.2. Note that the thermodynamic superstructure (Eq. 15) is not applied in this problem as the process is isothermal. A four-node hidden layer is used, with all activations comprised of three-point grids. The multiplication hyperparameter in the LeanKAN layer is set to 2 as per the recommendation of Ref. [33]. The training and testing data are sampled from the ground truth at a sparse time interval of 1 s, leading to 30 total data points for each individual species. The ChemKAN is trained on this data for epochs. Figure 3 shows example time-history species profiles for an (unseen) test case with the ground truth, noisy test data, and prediction by the learned ChemKAN model. With clean, noise-free data, the ChemKAN successfully learns the underlying model, showing good generalizability with accurate predictions at the training time steps, and smooth profiles in between.
@@
A common trait of experimental data from which machine learning techniques are tasked with inferring models is experimental uncertainty or noise, which is another avenue in which even properly parameterized deep learning techniques may overfit as they struggle to distinguish between genuine underlying trends and experimental noise. In fact, this is a known problem with standard KAN networks, where it was shown in Ref. [49] that one of the original cases in Liu’s original KAN work [23] fails with a small amount of added noise, a phenomenon that was further studied in more quantitative detail in Ref. [30]. In the current work, we further probe whether coupling to inherently noise-robust ODE solvers helps ChemKANs to extract useful models from increasingly noisy data. To do so, we task ChemKANs with extracting models from the same dataset with varying amounts of noise added (up to 15%) as shown in Fig. 3. Surprisingly, the ChemKAN, even with significant amounts of noise, demonstrates strong robustness and a capability to infer smooth and accurate solution profiles that correspond to the underlying true data. We provide detailed discussions on how ChemKAN performs well with comparison against DeepONet in the following subsections.
3.1.2 Neural scaling with noise-free data

### v2

@@ II.3.1 Vanilla KAN-ODEs
where each is a unique learnable activation function connecting a single input to a single output (thus ). In other words, each input is connected to each output with a unique learnable activation function (much like in an MLP, where each input is connected to each output with a unique learnable weight), leading to a total of activation functions connecting the and layers. We use RBF basis functions in the current work as was shown previously for KANs [34] and KAN-ODEs [29, 35], although the choice of is is flexible, and many other options have been proposed in the literature including B-splines [23], ReLU functions [36], and various other combinations [37, 38]. The AddKAN structure has an inherent problem of expression using only additive operations, limiting its concise expressivity for problems involving substantial use of the multiplication operator. Therefore, recent studies have proposed new layer structures to address this issue and improve parameter efficiency [39, 35]. Here we use LeanKAN [35], which has shown promise to be the most effective and efficient for both additive and multiplicative operations. The LeanKAN structure is achieved by summation of two separate terms, and , such that
where is the multiplication hyperparameter that dictates the number of multiplication input nodes for layer , and are the same univariate activation functions used in AddKAN (to be defined below). Overall, the LeanKAN formulation takes the matrix of activated inputs, computes sums and products in parallel according to to construct the intermediate , and finally sums the multiplication and addition components of to arrive at the next layer. Further derivation and implementation details are available for KANs [23], KAN-ODEs [29], and LeanKAN layers [35].
@@
where we define as the grid size, or number of superimposed basis functions used to construct a single activation. and are the learnable network parameters that make up , where scales each gridpoint’s RBF basis functions within the sum, and scales a single base activation function . and denote the input-output pair for which the current activation function is defined (e.g., the activation function with subscripts and connects the second input node to the third output node). The grid itself is defined by its individual gridpoints and gridpoint spacing (or RBF spreading parameter) . While Liu et al. [23] suggest periodically re-gridding these basis functions to ensure they are learning on the proper input range, here we instead normalize at each layer input using the hyperbolic tangent function as in more recent works [40, 41, 29, 35] for computational efficiency. The single base activation term in each is a Swish activation function [42].
II.3.2 Novel chemistry KAN (ChemKAN) architecture
@@
We define the terms in these equations line by line throughout this paragraph. In Eq. 13, we encode prior knowledge of the true species production/consumption rate relationship of Eq. 1 into the ChemKAN by computing the species-only production rate from the entire state input , via a KAN network (KAN) parameterized by . Then, from the true energy equation (Eq. 2), we recognize that the temperature rate is a simple linear combination of the species production rates, with scaling factors defined by the enthalpy and specific heat values (or alternatively, in Eq. 14 by the scalars in the linear mapping parameterized by ). Thus, the crux of the thermodynamic superstructure is a computationally trivial, simple linear sum of the already-evaluated outputs of the kinetic core, as shown in Eqs. 14 and 15. One level deeper, we recognize that a secondary effect in the true Eq. 2 is the dependence of the thermophysical parameters, specifically , on the temperature and species mixture. The error stemming from the first-order Linear() approximation’s failure to account for such thermophysical parameter variation is reflected in the term of Eq. 14, which in the final formulation of Eq. 15 is accounted for via a supplemental single-layer, single-output KAN correction carrying forward a functional dependence on the species and temperature inputs, parameterized by . Overall, ChemKAN is composed of a kinetic core structure and a thermodynamic superstructure that strongly mimic the true governing equations, operate largely in series, and include full sharing of all reaction and species production information. This architecture allows for versatile and flexible modeling by turning “on” or “off” the energy equation for standalone, kinetic core-only modeling or combined kinetic and thermodynamic modeling.
II.3.3 Network structure details
The ChemKAN structure proposed here is combined with the KAN-ODE framework detailed in Sec. II.3.1 to leverage both the physics-guided ChemKAN architecture and the proven dynamical system benefits offered by the KAN-ODE integration. For nomenclature consistency, we henceforth refer to both the newly designed network structure as well as its complete integration with an ODE solver as ChemKAN. For the relatively larger kinetic core responsible for learning all reaction information, we define the internal layer structure as
a two-layer structure per the notation introduced in Eq. 6. The first is an AddKAN layer [23] (which is equivalently [35] a LeanKAN layer with ). The second is a LeanKAN layer with to inject the multiplication operator. Further discussion of this specific two-layer form and its merits is available separately [35]. The thermodynamic correction structure simply comprises a scalar sum of univariate activations evaluated on the components of ,
@@ II.3.4 Loss function
Here denotes the thermochemical state normalized to the [0, 1] window by subtracting the minimum then dividing by the range. is the network prediction for this state quantity at time with the network parameters (including the kinetic, thermo, and correction parameters), while is the corresponding training data. is the number of datapoints in the temporal profiles. In the first MSE term, when training the kinetic core, as only species profiles are learned. For the thermodynamic superstructure, is used to train the added temperature output. We also provide an optional element conservation physics-informed loss term, or PINN term [43, 44], to encourage the ChemKAN to find models that obey physical laws. There, is the number of elements in the data (i.e., H, O, N). For a given element , the element conservation term begins by computing the mass fraction difference across all species between the current ChemKAN timestep and the initial condition. This is then converted to an elemental conservation difference via , the atom count of element in species ; , the atomic mass of element ; and , the molar mass of species . This conserved term is computed at all timesteps across all elements , and then weighted by (here ). In the examples below, we use only the MSE loss for all biodiesel model inference results, and the MSE loss with the PINN term for H2 model acceleration.
II.3.5 Training Process
We highlight again that this architecture incorporates significant inductive bias by separating the inference of the kinetic behavior and thermodynamic behavior into distinct layers and activations within the complete structure, giving the network an explicit functional coupling between the temperature rate and the species rates to ease its training burden while retaining an accurate model of the real exothermic and endothermic behavior of the system. In itself, this serially run inductive split should theoretically ease some of the nonlinear training difficulties observed with ChemNODE [17] while retaining a single cohesive and information-sharing structure. Additionally, and more importantly to stable training and convergence behavior, it allows us to separate the training itself into a kinetic and a thermodynamic stage to greatly reduce the inference burden, as discussed here:
Training stage 1—core kinetics : All inputs are used, and only the species production rate outputs are learned (see Eq. 13). The thermodynamic superstructure of Eqs. 14 and 15 is not used in this stage, and the input temperatures are simply read in from the training data to provide the kinetic core of the network with a simpler training task. See the grey highlight in Fig. 2. For cases without heat release, this step in isolation is sufficient for a complete model. With heat release, we move to stage 2 once stage 1 is converged.
@@ II.3.5 Training Process
This two-stage training process contrasts with the stage training process of ChemNODE [17], which requires networks to all learn their own distinct representations of what we know to be shared kinetic and thermodynamic governing laws. The current ChemKAN approach, while trained in two distinct stages, has an overall structure resembling a simple feedforward KAN thanks to its stacked design, where all kinetic and thermodynamic information is shared across all inputs and outputs. For downstream evaluation, a single forward pass through the combined network of Fig. 2, or alternatively through Eqs. 1315, predicts the complete thermochemical state vector .
While adjoint sensitivity analysis was used in the original KAN-ODE paper [29] to compute the gradients of a loss function , in the current work we note the numerical instability present in many chemical kinetic modeling problems due to numerical stiffness [45, 22, 18, 17, 3], potentially leading to failures with adjoint sensitivity analysis in the training process [3]. To prevent this, we implement forward sensitivity analysis to mitigate this potential stiffness issue in the ODE solver [3]. ChemKANs and their corresponding forward sensitivity equations are solved using an ODE integrator of Tsit5 (Tsitouras 5/4 Runge-Kutta method [46]). The learnable parameters are trained by the Adam optimizer [47] with a learning rate of .
II.4 Data generation for case studies
We introduce two chemical reaction systems to demonstrate ChemKANs as a multi-purpose modeling technique. The first case of biodiesel production will illustrate the effectiveness of ChemKANs for inferring a chemical kinetic model purely from noisy data. Then, we show computational acceleration for high-fidelity simulations using ChemKANs in a hydrogen-air combustion example subject to significant numerical stiffness.
II.4.1 Model inference from noisy data – biodiesel production
@@
where the reaction rates scale with temperature via the standard Arrhenius law as , with for the three reactions. As in Ref. [5], we generate data in this case using kcal/mol and , with isothermal experiments at temperatures randomly sampled in the range of K to K. We define the species scalar quantities here as concentrations rather than mass fractions, to match the convention with the governing equations. Initial TG and ROH concentrations are randomly sampled uniformly between and with all other intermediate and output species initialized at zero. 20 training data sets and 10 testing data sets are generated, with 30-second time windows in both consisting of sampled points. The temperature-dependent yet isothermal reaction rates present in this system motivate the use of the kinetic ChemKAN core structure only (see the Hydrogen example below for use of the kinetic core together with the thermodynamic superstructure).
@@
In this case we additionally probe the effectiveness of ChemKANs in the presence of significant experimental noise. To do so, we add increasing amounts of noise to the data and evaluate ChemKAN’s performance against that of a standard Deep Operator Network (DeepONet) [20].
II.4.2 Model acceleration – hydrogen-air combustion
@@
Despite its relatively simple chemistry, hydrogen-air combustion is subject to numerical stiffness, leading to high computational costs and numerical instability in the ODE solver. The main goal of the ChemKAN here is to achieve high accuracy in species and temperature profile reconstruction while reducing computational cost compared to the detailed chemistry solver. We generate training data in Cantera [50], using the H2/O2 mechanism from GRI-Mech 3.0 [51] (9 species, 29 reactions). Data are generated at all 36 combinations of initial temperatures in {950, 1000, 1050, 1100, 1150, 1200} K and equivalence ratios in {0.7, 0.9, 1.1, 1.3, 1.5}, with the [1150 K, 1.3] case withheld as unseen testing data. The problem setup here is largely identical to that studied originally in ChemNODE [17] to facilitate fair comparison, although here we add the single withheld testing dataset to probe ChemKAN’s robustness (thus 35 training datasets are used, compared to the 36 in ChemNODE).
@@
In addition to significantly increased numerical stiffness and complexity, a key fundamental difference between the current combustion system and the previous biodiesel synthesis case is the presence of substantial two-way temperature coupling as per Eqs. 12. To account for this, we include the thermodynamic ChemKAN superstructure and the two-stage training process outlined in Secs. II.3.2II.3.5.
III Results and Discussion
III.1 Biodiesel model inference from noisy data
@@
We begin with an analysis of the biodiesel model inference of the ChemKAN (the kinetic core only for this isothermal case), and a comparison against a traditional DeepONet for an identical task. DeepONets were selected as the target of comparison due to their general use in the scientific machine learning community, to provide a baseline against which to evaluate our first ChemKAN results. Note that comparison against ChemNODE is saved for the hydrogen-air combustion case in Sec. III.2.
III.1.1 Model performance of the proposed ChemKAN architecture
@@
The kinetic core (Eq. 13) of the ChemKAN comprises two layers: one AddKAN and one LeanKAN, as discussed in Sec. II.3.2. Note that the thermodynamic superstructure (Eq. 15) is not applied in this problem as the process is isothermal. A four-node hidden layer is used, with all activations comprised of three-point grids. The multiplication hyperparameter in the LeanKAN layer is set to 2 as per the standard LeanKAN formulation [35]. The training and testing data are sampled from the ground truth at a sparse time interval of 1 s, leading to 30 total data points for each individual species. The ChemKAN is trained on this data for epochs. The leftmost column of Figure 3 shows example time-history species profiles for an unseen test case with the ground truth, unseen test data, and prediction by the learned ChemKAN model. With clean, noise-free data, the ChemKAN successfully learns the underlying model and shows good generalizability, with accurate predictions at the training time steps and smooth profiles in between.
@@
A common characteristic of experimental data used for machine learning model inference is the presence of uncertainty or noise, which can cause even well-parameterized deep learning models to overfit as they struggle to distinguish genuine underlying trends from experimental artifacts. This is no different in traditional KANs, where small amounts of noise have been shown to severely limit inference capabilities [32]. In the current work, we further probe whether coupling to inherently noise-robust ODE solvers helps ChemKANs to extract useful models from increasingly noisy data. To do so, we task ChemKANs with extracting models from the same dataset with varying amounts of noise added (up to 15%) as shown in Fig. 3. Surprisingly, the ChemKAN, even with significant amounts of noise, demonstrates strong robustness and a capability to infer smooth and accurate solution profiles that correspond well to the underlying true data. We provide detailed discussions on how ChemKAN performs compared to DeepONet in the following subsections.
III.1.2 Neural scaling with noise-free data